OpenStudy (anonymous):

if the measurement of angle AFB= 8x-6 and the measurement of BFC is 14x+8 , find the value of x so that angle AFC is the right angle

OpenStudy (anonymous):

8x -6 = AFB 14x + 8 = BFC AFC is 90 AFB + BFC = 90 8x -6 + 14x +8 = 90 22x +2 = 90 22x =88 x = 44

OpenStudy (anonymous):

wouldnt it be x=4